# spacing=nonfrench

### v2

@@ C.3 Chemical resistance
$R_{i}$ $=\frac{\alpha_{i}}{\varphi_{i}}$ (25aibaa)
$R_{ra}$ $=\frac{2(1+\varepsilon\theta_{g})}{\kappa^{\prime}\rho}$ (25aibab)
$R_{rb}$ $=\frac{2(1-\varepsilon\theta_{g})}{\kappa^{\prime}\rho}$ (25aibac)
$R_{d1}$ $=\frac{2\kappa\lambda}{\kappa^{\prime}\delta}$ (25aibad)
$R_{d2}$ $=\frac{2}{\kappa^{\prime}\delta\kappa\lambda}$ (25aibae)
$R_{r}$ $=\frac{4}{\kappa^{\prime}\rho}$ (25aibaf)
$R_{d}$ $=\frac{4}{\kappa^{\prime}\lambda^{\prime}\delta}$ (25aibag)
$R_{\text{tot}}$ $=\frac{4}{\kappa^{\prime}}\left(\frac{1}{\rho}+\frac{1}{\lambda^{\prime}\delta}\right)$ (25aibah)
@@ C.4 Entropy yield
The ratio $y_{k}=\nicefrac{{\sigma_{k}}}{{\sigma}}$ provides the fraction of entropy dissipated through the process $k$. This can be linked to the energy yield, which is the fraction of the energy dissipated through a specific process of interest. This can be calculated as:
$y_{k}$ $=\frac{\sigma_{k}}{\sigma}$ (25aibba)
$=\frac{R_{k}\varphi^{2}}{R_{\text{tot}}\varphi^{2}}=\frac{R_{k}}{R_{\text{tot}}}$ (25aibbb)
@@
$y_{r}$ $=\frac{\lambda^{\prime}\delta}{\lambda^{\prime}\delta+\rho}$ (25aibca)
$y_{d}$ $=\frac{\rho}{\lambda^{\prime}\delta+\rho}$ (25aibcb)
C.5 General expression
All $\varphi$, $\alpha_{k}$ and $\sigma_{k}$ values can be directly derived from the expression of $R_{k}$ in Eqs. (25aiba) and $\alpha$ from Eq. (25aias):
$\varphi$ $=\frac{\alpha}{R_{\text{tot}}}$ (25aibda)
$\alpha_{k}$ $=\frac{R_{k}}{R_{\text{tot}}}\alpha$ (25aibdb)
$\sigma_{k}$ $=\frac{R_{k}}{R_{\text{tot}}^{2}}\alpha^{2}$ (25aibdc)
@@ C.5 General expression
$\varphi$ $=\frac{\lambda^{\prime}\delta\rho}{\lambda^{\prime}\delta+\rho}\varphi_{0}$ (25aibea)
@@
$\varphi_{0}$ $=\frac{1}{2}\kappa^{\prime}\eta\theta_{g}$ (25aibeb)
@@
$\alpha_{r}$ $=\frac{\lambda^{\prime}\delta}{\lambda^{\prime}\delta+\rho}\alpha_{0}$ (25aibec)
$\alpha_{d}$ $=\frac{\rho}{\lambda^{\prime}\delta+\rho}\alpha_{0}$ (25aibed)
@@
$\alpha_{0}$ $=2\eta\theta_{g}$ (25aibee)
@@
$\sigma_{r}$ $=\frac{(\lambda^{\prime}\delta)^{2}\rho}{(\lambda^{\prime}\delta+\rho)^{2}}\sigma_{0}$ (25aibef)
$\sigma_{d}$ $=\frac{\lambda^{\prime}\delta\rho^{2}}{(\lambda^{\prime}\delta+\rho)^{2}}\sigma_{0}$ (25aibeg)
@@
$\sigma_{0}$ $=\kappa^{\prime}(\eta\theta_{g})^{2}$ (25aibeh)
for the entropy dissipation.
This results in a reference steady state ($[\alpha_{0},\varphi_{0},\sigma_{0}]$), which depends on the thermokinetic characteristics of the chemical reaction. The state of a given system can be retrieved from this reference state based on Eqs. (25aibe).
C.6 Entropy dissipation variations
@@
$\left(\frac{\partial\sigma_{r}}{\partial\rho}\right)_{\delta,\sigma_{0}}$ $=\frac{(\lambda^{\prime}\delta)^{2}}{\left(\delta\lambda^{\prime}+\rho\right)^{3}}\sigma_{0}\left(\lambda^{\prime}\delta-\rho\right)$ (25aibfa)
$\left(\frac{\partial\sigma_{r}}{\partial\delta}\right)_{\rho,\sigma_{0}}$ $=\frac{2\lambda^{\prime}\delta\rho^{2}}{\left(\delta\lambda^{\prime}+\rho\right)^{3}}\sigma_{0}$ (25aibfb)
@@
$\left(\frac{\partial\sigma_{d}}{\partial\rho}\right)_{\delta,\sigma_{0}}$ $=\frac{2(\lambda^{\prime}\delta)^{2}\rho}{\left(\delta\lambda^{\prime}+\rho\right)^{3}}\sigma_{0}$ (25aibga)
$\left(\frac{\partial\sigma_{d}}{\partial\delta}\right)_{\rho,\sigma_{0}}$ $=\frac{\rho^{2}}{\left(\delta\lambda^{\prime}+\rho\right)^{3}}\sigma_{0}\left(\rho-\lambda^{\prime}\delta\right)$ (25aibgb)
@@ D.1 Expression of the PDE model
$\frac{\partial u_{1}}{\partial\bar{t}}$ $=d_{1}\frac{\partial^{2}u_{1}}{\partial\bar{x}^{2}}-\varphi_{r}(\bar{x}),$ (25aibha)
$\frac{\partial u_{2}}{\partial\bar{t}}$ $=d_{2}\frac{\partial^{2}u_{2}}{\partial\bar{x}^{2}}+\varphi_{r}(\bar{x}),$ (25aibhb)
@@
Chemical transport is now defined as molecular diffusion; it cannot be treated anymore as a chemical reaction. The rates of entropy production of the chemical reaction ($\sigma_{r}$) and chemical diffusion of compound $U_{i}$ ($\sigma_{d,i}$) are expressed as[46]:
$\sigma_{r}$ $=\varphi_{r}\alpha_{r}=(\bar{k}_{+}u_{1}-\bar{k}_{-}u_{2})\ln\frac{\bar{k}_{+}u_{1}}{\bar{k}_{-}u_{2}}$ (25aibia)
$\sigma_{d,i}$ $=\frac{d_{i}}{u_{i}}\left(\frac{\partial u_{i}}{\partial\bar{x}}\right)^{2}$ (25aibib)
@@
The steady-state conditions must consider the spatial extension of the system. In steady state, the total concentrations of each compound $U_{1}$ and $U_{2}$ are constant. This implies that the global reaction is zero, i.e.:
@@
Because $\sigma$ is necessarily a positive value, its integral value is calculated over the full system extension:

### v3

@@ C.3 Chemical resistance
$R_{i}$ $=\frac{\alpha_{i}}{\varphi_{i}}$ (4bvbwcra)
$R_{ra}$ $=\frac{2(1+\varepsilon\theta_{g})}{\kappa^{\prime}\rho}$ (4bvbwcrb)
$R_{rb}$ $=\frac{2(1-\varepsilon\theta_{g})}{\kappa^{\prime}\rho}$ (4bvbwcrc)
$R_{d1}$ $=\frac{2\kappa\lambda}{\kappa^{\prime}\delta}$ (4bvbwcrd)
$R_{d2}$ $=\frac{2}{\kappa^{\prime}\delta\kappa\lambda}$ (4bvbwcre)
$R_{r}$ $=\frac{4}{\kappa^{\prime}\rho}$ (4bvbwcrf)
$R_{d}$ $=\frac{4}{\kappa^{\prime}\lambda^{\prime}\delta}$ (4bvbwcrg)
$R_{\text{tot}}$ $=\frac{4}{\kappa^{\prime}}\left(\frac{1}{\rho}+\frac{1}{\lambda^{\prime}\delta}\right)$ (4bvbwcrh)
@@ C.4 Entropy yield
The ratio $y_{k}=\nicefrac{{\sigma_{k}}}{{\sigma}}$ provides the fraction of entropy produced by the process $k$. This can be linked to the energy yield, which is the fraction of the energy produced by a specific process of interest. This can be calculated as:
$y_{k}$ $=\frac{\sigma_{k}}{\sigma}$ (4bvbwcsa)
$=\frac{R_{k}\varphi^{2}}{R_{\text{tot}}\varphi^{2}}=\frac{R_{k}}{R_{\text{tot}}}$ (4bvbwcsb)
@@
$y_{r}$ $=\frac{\lambda^{\prime}\delta}{\lambda^{\prime}\delta+\rho}$ (4bvbwcta)
$y_{d}$ $=\frac{\rho}{\lambda^{\prime}\delta+\rho}$ (4bvbwctb)
C.5 System characteristics
All $\varphi$, $\alpha_{k}$ and $\sigma_{k}$ values can be directly derived from the expression of $R_{k}$ in Eqs. (4bvbwcr) and $\alpha$ from Eq. (4bvbwcj):
$\varphi$ $=\frac{\alpha}{R_{\text{tot}}}$ (4bvbwcua)
$\alpha_{k}$ $=\frac{R_{k}}{R_{\text{tot}}}\alpha$ (4bvbwcub)
$\sigma_{k}$ $=\frac{R_{k}}{R_{\text{tot}}^{2}}\alpha^{2}$ (4bvbwcuc)
@@ C.5 System characteristics
$\varphi$ $=\frac{\lambda^{\prime}\delta\rho}{\lambda^{\prime}\delta+\rho}\varphi_{0}$ (4bvbwcva)
@@
$\varphi_{0}$ $=\frac{1}{2}\kappa^{\prime}\eta\theta_{g}$ (4bvbwcvb)
@@
$\alpha_{r}$ $=\frac{\lambda^{\prime}\delta}{\lambda^{\prime}\delta+\rho}\alpha_{0}$ (4bvbwcvc)
$\alpha_{d}$ $=\frac{\rho}{\lambda^{\prime}\delta+\rho}\alpha_{0}$ (4bvbwcvd)
@@
$\alpha_{0}$ $=2\eta\theta_{g}$ (4bvbwcve)
@@
$\sigma_{r}$ $=\frac{(\lambda^{\prime}\delta)^{2}\rho}{(\lambda^{\prime}\delta+\rho)^{2}}\sigma_{0}$ (4bvbwcvf)
$\sigma_{d}$ $=\frac{\lambda^{\prime}\delta\rho^{2}}{(\lambda^{\prime}\delta+\rho)^{2}}\sigma_{0}$ (4bvbwcvg)
$\sigma=\sigma_{r}+\sigma_{d}$ $=\frac{\lambda^{\prime}\delta\rho}{\lambda^{\prime}\delta+\rho}\kappa^{\prime}\cdot\eta^{2}\theta_{g}^{2}$ (4bvbwcvi)
@@
$\sigma_{0}$ $=\kappa^{\prime}(\eta\theta_{g})^{2}$ (4bvbwcvj)
for the entropy production.
This results in a reference steady state ($[\alpha_{0},\varphi_{0},\sigma_{0}]$), which depends on the thermokinetic characteristics of the chemical reaction. The state of a given system can be retrieved from this reference state based on Eqs. (4bvbwcv).
C.6 Entropy production variations
@@
$\left(\frac{\partial\sigma_{r}}{\partial\rho}\right)_{\delta,\sigma_{0}}$ $=\frac{(\lambda^{\prime}\delta)^{2}}{\left(\delta\lambda^{\prime}+\rho\right)^{3}}\sigma_{0}\left(\lambda^{\prime}\delta-\rho\right)$ (4bvbwcwa)
$\left(\frac{\partial\sigma_{r}}{\partial\delta}\right)_{\rho,\sigma_{0}}$ $=\frac{2\lambda^{\prime}\delta\rho^{2}}{\left(\delta\lambda^{\prime}+\rho\right)^{3}}\sigma_{0}$ (4bvbwcwb)
@@
$\left(\frac{\partial\sigma_{d}}{\partial\rho}\right)_{\delta,\sigma_{0}}$ $=\frac{2(\lambda^{\prime}\delta)^{2}\rho}{\left(\delta\lambda^{\prime}+\rho\right)^{3}}\sigma_{0}$ (4bvbwcxa)
$\left(\frac{\partial\sigma_{d}}{\partial\delta}\right)_{\rho,\sigma_{0}}$ $=\frac{\rho^{2}}{\left(\delta\lambda^{\prime}+\rho\right)^{3}}\sigma_{0}\left(\rho-\lambda^{\prime}\delta\right)$ (4bvbwcxb)
@@ D.1 Expression of the PDE model
$\frac{\partial u_{1}}{\partial\bar{t}}$ $=d_{1}\frac{\partial^{2}u_{1}}{\partial\bar{x}^{2}}-\varphi_{r}(\bar{x}),$ (4bvbwcya)
$\frac{\partial u_{2}}{\partial\bar{t}}$ $=d_{2}\frac{\partial^{2}u_{2}}{\partial\bar{x}^{2}}+\varphi_{r}(\bar{x}),$ (4bvbwcyb)
@@
Chemical transport is now defined as molecular diffusion; it cannot be treated anymore as a chemical reaction. The rates of entropy production of the chemical reaction ($\sigma_{r}$) and chemical diffusion of compound $U_{i}$ ($\sigma_{d,i}$) are expressed as [36, 37]:
$\sigma_{r}=\varphi_{r}\cdot\alpha_{r}$ $=(\bar{k}_{+}u_{1}-\bar{k}_{-}u_{2})\cdot\ln\frac{\bar{k}_{+}u_{1}}{\bar{k}_{-}u_{2}}$ (4bvbwcza)
$\sigma_{d,i}=\varphi_{d}\cdot\alpha_{d}$ $=j_{i}\cdot\frac{1}{u_{i}}\left(\frac{\partial u_{i}}{\partial\bar{x}}\right)$ (4bvbwczb)
$=\frac{d_{i}}{u_{i}}\left(\frac{\partial u_{i}}{\partial\bar{x}}\right)^{2}$ (4bvbwczc)
@@ D.2 Steady state characteristics
The steady-state conditions must consider the spatial extension of the system. In steady state, the total concentrations of each compound $U_{1}$ and $U_{2}$ are constant. This implies that the global reaction flux is zero, i.e.:
@@
A positive rate of entropy production $\sigma$ is present at each system location. The full entropy production comes down to the integral value is calculated over the full system extension: